# Training robust and generalizable quantum models

### v2

@@ A Lipschitz bound of a model $f$ is any $L>0$ satisfying
These insights confirm the observation by [41, 42] that fixed-encoding quantum models are neither sensitive to data perturbations nor to overfitting. On the one hand, resilience against these two phenomena is a desirable property. However, the above discussion also implies that Lipschitz bound regularization, which is a systematic and effective tool for influencing robustness and generalization [2, 5, 11, 12], cannot be implemented for fixed-encoding quantum models to improve robustness and generalization. Indeed, our robustness simulations in Figure 1 show that the fixed-encoding model has a considerably higher Lipschitz bound than the considered trainable-encoding models. This implies a significantly worse robustness w.r.t. data perturbations and, therefore, leads to a rapidly decreasing test accuracy for larger noise levels. Further, regularizing the parameters $\varphi$ as suggested, e.g., by [43], does not affect the Lipschitz bound and, therefore, cannot be used to improve the robustness, see the supplementary material (Section III) for a numerical demonstration.
Conclusion.—In this letter, we study robustness and generalization properties of quantum models based on Lipschitz bounds. We derive Lipschitz bounds based on the norm of the data encoding which not only suggest a novel regularization strategy to improve robustness and generalization, but also highlight the role of trainable encodings. The numerical results confirm our theoretical findings, showing the existence of a sweet spot for the regularization parameter for which our training scheme improves both robustness as well as generalization compared to a non-regularized training scheme. We expect manifold applications of Lipschitz bounds in QML, ranging from concrete next steps (e.g., nonlinear encodings based on Lipschitz-bounded classical neural networks [2, 5, 7, 8, 9, 10]) to broader perspectives for transferring classical ML concepts to QML.
@@
Code Availability.—The source code for the numerical case studies, as well as the data generation and analysis scripts are publicly accessible on GitHub via https://github.com/daniel-fink-de/training-robust-and-generalizable-quantum-models.
@@
Acknowledgment.—This work was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 - 390740016. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). This work was also supported by the German Federal Ministry of Economic Affairs and Climate Action through the project AutoQML (grant no. 01MQ22002A).
Supplementary material
The supplementary material contains additional technical details on the theoretical analysis (formal statements and proofs) as well as the numerical results (numerical setup and additional results).
I. Lipschitz bounds of quantum models
@@
In this section, we study Lipschitz bounds of the quantum models considered in the main manuscript. We first derive a Lipschitz bound which is less tight than the one in (5) but can be shown using a simple concatenation argument (Section I.A). Next, in Section I.B, we prove that (5) is indeed a Lipschitz bound.
A. Simple Lipschitz bound based on concatenation
@@
Theorem I.1. The following is a Lipschitz bound of $f_{\Theta}$:
@@
$\lVert f(x)-f(y)\rVert$ $\leq L_{1}\lVert f_{2}\circ\dots\circ f_{h}(x)-f_{2}\circ\dots\circ f_{h}(y)\rVert$
$\leq\dots\leq L_{1}L_{2}\cdots L_{h}\lVert x-y\rVert.$ (14)
@@
Lipschitz bound of $g_{\mathrm{unitary}}$: It follows from [21, Theorem 2.2] that $L_{\mathrm{unitary}}=\sum_{j=1}^{N}\lVert H_{j}\rVert$ is a Lipschitz bound of $g_{\mathrm{unitary}}$.
@@
B. Proof that (5) is a Lipschitz bound
@@
$\mathrm{d}U_{\Theta}(x)=$ $(\mathrm{d}U_{N,\Theta_{N}}(x))U_{N-1,\Theta_{N-1}}(x)\cdots U_{1,\Theta_{1}}(x)$ (22)
@@
$\mathrm{d}U_{\Theta}(x)(v)=-i\left(H_{N}U_{\Theta}(x)w_{N}^{\top}v+\dots+U_{\Theta}(x)H_{1}w_{1}^{\top}v\right).$ (27)
@@
$J_{\Theta}(x)=-i\left(H_{N}U_{\Theta}(x)\ket{0}w_{N}^{\top}+\dots+U_{\Theta}(x)H_{1}\ket{0}w_{1}^{\top}\right).$ (28)
@@
Thus, $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ is a Lipschitz bound of $U_{\Theta}(x)\ket{0}$ [44, p. 356].
@@
II. Full version and proof of Theorem 1
We first state the main result in a general supervised learning setup, before applying it to the quantum model considered in the main manuscript. Consider a supervised learning setup with data samples $\{x_{k},y_{k}\}_{k=1}^{n}$ drawn independently and identically distributed from $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ according to some probability distribution $P$. We define the $\epsilon$-covering number of $\mathcal{Z}$ as follows.
Definition II.1. (adapted from [6, Definition 1]) We say that $\hat{\mathcal{Z}}$ is an $\epsilon$-cover of $\mathcal{Z}$, if, for all $z\in\mathcal{Z}$, there exists $\hat{z}\in\hat{\mathcal{Z}}$ such that $\lVert z-\hat{z}\rVert\leq\epsilon$. The $\epsilon$-covering number of $\mathcal{Z}$ is
@@ II. Full version and proof of Theorem 1
    $\mathcal{Z}$ is compact such that $M\coloneqq\sup_{y,y^{\prime}\in\mathcal{Y}}\ell(y,y^{\prime})$ is finite, and
@@
For the following proof, we invoke the concept of $(K,\epsilon)$-robustness (adapted from [6, Definition 2]): The classifier $f$ is $(K,\epsilon)$-robust for $K\in\mathbb{N}$ and $\epsilon\in\mathbb{R}$, if $\mathcal{Z}$ can be partitioned into $K$ disjoint sets, denoted by $\{C_{i}\}_{i=1}^{K}$, such that the following holds: For all $k=1,\dots,n$, $(x,y)\in\mathcal{Z}$, $i=1,\dots,K$, if $(x_{k},y_{k}),(x,y)\in C_{i}$, then
@@
where we use the Lipschitz bound $L_{\ell}$ of $\ell$, the triangle inequality, and the Lipschitz bound $L_{f}$ of $f$, respectively. Using [6, Theorem 6], we infer that $f$ is $(\mathcal{N}(\frac{\gamma}{2},\mathcal{Z}),L_{\ell}\max\{1,L_{f}\}\gamma)$-robust for all $\gamma>0$. It now follows from [6, Theorem 1] that, for any $\delta>0$, with probability at least $1-\delta$ inequality (33) holds. ∎
Let us now combine the Lipschitz bound (5) and Lemma II.1 to state a tailored generalization bound for the considered class of quantum models $f_{\Theta}$, thus proving Theorem 1 in the main text.
Theorem II.1. Suppose
@@
    $\mathcal{Z}$ is compact such that $M\coloneqq\sup_{y,y^{\prime}\in\mathcal{Y}}\ell(y,y^{\prime})$ is finite.
@@
$|\mathcal{R}(f)-\mathcal{R}_{n}(f)|\leq$ $\gamma L_{\ell}\max\Big{\{}1,2\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert\Big{\}}$
$+M\sqrt{\frac{2\mathcal{N}(\frac{\gamma}{2},\mathcal{Z})\ln 2+2\ln(\frac{1}{\delta})}{n}}.$ (36)
Theorem II.1 shows that the size of the data encoding and of the observable directly influences the generalization performance of the quantum model $f_{\Theta}$. In particular, for smaller values of $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and $\lVert\mathcal{M}\rVert$, the expected loss is closer to the empirical loss. The right-hand side of (36) contains two terms: The first one depends on the parameters of the quantum model $f_{\Theta}$ and characterizes its robustness via the derived Lipschitz bound, whereas the second term decays with increasing data length $n$. While (33) holds for arbitrary values of $\gamma>0$, it is not immediate which value of $\gamma$ leads to the smallest possible bound: Smaller values of $\gamma$ decrease the first term but increase $\mathcal{N}(\frac{\gamma}{2},\mathcal{Z})$ in the second term (and vice versa).
In contrast to existing QML generalization bounds [30, 31, 32, 33, 34], Theorem II.1 explicitly highlights the role of the Lipschitz bound. Using the additional flexibility of the parameter $\gamma$, it can be shown that the generalization bound (36) converges to zero when the data length $n$ approaches infinity. To this end, we use [45, Lemma 6.27] to upper bound the covering number
@@
$|\mathcal{R}(f)-\mathcal{R}_{n}(f)|\leq$ $\frac{1}{n^{\frac{1}{2D+2}}}L_{\ell}\max\Big{\{}1,2\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert\Big{\}}$
$+M\sqrt{\frac{2\ln 2\left(6R\right)^{D+1}}{\sqrt{n}}+\frac{2\ln(\frac{1}{\delta})}{n}},$ (38)
@@
III. Numerics: setup and further results
In the following, we provide details regarding the setup of our numerical results (Section .1) and we present further numerical results regarding parameter regularization in quantum models with fixed encoding (Section .2).
.1 A. Numerical setup
@@ III. Numerics: setup and further results
All numerical simulations within this work were performed using the Python QML library PennyLane [46]. As device, we used the noiseless simulator ”lightning.qubit” together with the adjoint differentiation method, to enable fast and memory efficient gradient computations. In order to solve the optimization problem in (8), we apply the ADAM optimizer using a learning rate of $\eta=0.1$ and the suggested values for all other hyperparameters [47]. Furthermore, we run $200$ epochs throughout and train $12$ models based on different initial parameters for varying regularization parameters $\lambda\geq 0$. Adding the regularization does not introduce significant computational overhead as the evaluation of the cost only involves a weighted sum of the terms $\|w_{j}\|^{2}$. As final model, we take the set of parameters for the model with minimal cost over all runs and epochs. Furthermore, the training as well as the robustness and generalization analysis were parallelized using Dask [48].
@@
$U_{\mathrm{Rot}}(\alpha,\beta,\gamma)$ $=R_{Z}(\gamma)R_{Y}(\beta)R_{Z}(\alpha)\,,$ (42)
$=\begin{pmatrix}e^{-\frac{i}{2}(\gamma+\alpha)}\cos(\frac{\beta}{2})&-e^{\frac{i}{2}(\gamma-\alpha)}\sin(\frac{\beta}{2})\\ e^{-\frac{i}{2}(\gamma-\alpha)}\sin(\frac{\beta}{2})&e^{\frac{i}{2}(\gamma+\alpha)}\cos(\frac{\beta}{2})\end{pmatrix}\,.$ (43)
Figure 4: Results for the generalization simulations for the circle classification problem when using the fixed-encoding quantum model (11) and regularization of the parameters $\varphi$. The training and test setup are identical to simulations shown in Figures 1 and 2. We plot the dependency of the test accuracy and the Lipschitz bound on the hyperparameter $\lambda$ entering the regularized training problem (44). Further, we plot the test accuracy for the best solution of the unregularized training problem.
@@
.2 B. Regularization in quantum models with fixed encoding
In the main text, we have seen that the Lipschitz bound of the quantum model with fixed encoding $f_{\varphi}^{\mathrm{f}}(x)$ in (11) cannot be adapted by changing the parameters $\varphi$. As a result, it is not possible to use Lipschitz bound regularization for improving robustness and generalization. In the following, we investigate whether regularization of $\varphi$ can instead be used to improve generalization performance. More precisely, we consider the same numerical setup as for our generalization results with trainable encoding depicted in Figure 2. The main difference is that we consider a fixed-encoding quantum model $f_{\varphi}^{\mathrm{f}}(x)$ (compare Figure 6) which is trained via the following regularized training problem
$\min_{\varphi}\>\>\frac{1}{n}\sum_{k=1}^{n}\ell(f_{\varphi}^{\mathrm{f}}(x_{k}),y_{k})+\lambda\sum_{j=1}^{L}\lVert\varphi_{j}\rVert^{2}.$ (44)
The regularization with hyperparameter $\lambda>0$ aims at keeping the norms of the angles $\varphi_{j}$ small. Figure 4 depicts the test accuracy and Lipschitz bound of the resulting quantum model for different regularization parameters $\lambda$. First, note that the Lipschitz bound is indeed constant for all choices of $\lambda$ due to the fixed encoding. Comparing Figures 2 and 4, we see that the trainable encoding yields a significantly higher test accuracy in comparison to the fixed encoding. Moreover, the influence of the regularization parameter on the test accuracy is much less pronounced for the fixed encoding than for the trainable encoding, confirming our previous discussion on the benefits of trainable encodings. The test accuracy is not entirely independent of $\lambda$ since 1) regularization of the parameters $\varphi_{j}$ influences the optimization and can improve or deteriorate convergence and 2) biasing $\varphi$ towards zero may be beneficial if the underlying ground-truth distribution is better approximated by a quantum model with small values $\varphi$. Whether 2) brings practical benefits is, however, highly problem-specific as it depends on the distribution generating the data and on the circuit ansatz.

### v3

@@ 5 Benefits of trainable encodings
These insights confirm the observation by [56, 57] that fixed-encoding quantum models are neither sensitive to data perturbations nor to overfitting. On the one hand, resilience against these two phenomena is a desirable property. However, the above discussion also implies that Lipschitz bound regularization, which is a systematic and effective tool for influencing robustness and generalization [2, 6, 16, 18], cannot be implemented for fixed-encoding quantum models to improve robustness and generalization. Indeed, our robustness simulations in Figure 2 show that the fixed-encoding model has a considerably higher Lipschitz bound than all the considered trainable-encoding models. This implies a significantly worse robustness w.r.t. data perturbations and, therefore, leads to a rapidly decreasing test accuracy for larger noise levels. Further, regularizing the parameters $\phi$ as suggested, e.g., by [58], does not affect the Lipschitz bound and, therefore, cannot be used to improve the robustness. In Appendix C, we study the effect of regularizing the $\phi_{j}$’s on generalization. We find that the influence of regularizing the $\phi_{j}$’s on the test accuracy is limited and likely dependent on the specific ground-truth distribution generating the data and the chosen circuit ansatz.
To conclude, our results show that training the encoding in quantum models not only increases the expressivity but also leads to superior robustness and generalization properties.
6 Conclusion
In this paper, we studied robustness and generalization properties of quantum models based on Lipschitz bounds. Lipschitz bounds are a well-established tool in the classical ML literature which not only quantify adversarial robustness but are also closely connected to generalization performance. We derived Lipschitz bounds based on the size of the data encoding which we then used to study robustness and generalization of quantum models. Given that our generalization bound explicitly involves the parameters of the data encoding, it does not face the limitations of uniform generalization bounds [49]. Further, our theoretical results highlight the role of trainable encodings combined with regularization techniques for obtaining robust and generalizable quantum models. The numerical results confirm our theoretical findings, showing the existence of a sweet spot for the regularization parameter for which our training scheme improves both robustness as well as generalization compared to a non-regularized training scheme. It is important to emphasize that these numerical results with specific choices of rotation and entangling gates are mainly used for illustration, but our theoretical framework applies to all quantum models that can be written as (4) and, therefore, also allow for different rotation gates, entangling layers, or even parametrized multi-qubit gates.
While our results indicate the potential of using Lipschitz bounds and regularization techniques in QML, it opens up various promising directions for future research. First and foremost, transferring existing research on Lipschitz bounds in classical ML to the QML setting provides a systematic framework for handling robustness and generalization, beyond the first results presented in this paper. For example, while we only consider quantum models with affine encodings $w_{j}^{\top}x+\theta_{j}$, it would be interesting to extend our results to more general, nonlinear encodings. Classical neural networks are ideal candidates for realizing a nonlinear encoding since their Lipschitz properties are well-studied [2, 6, 9, 10, 11, 12, 13, 14, 15], which would allow to train hybrid quantum-classical models which are not only expressive but also admit desirable robustness and generalization properties. Finally, although we focus on variational quantum models, the basic principles of our results are transferrable to different quantum models, including quantum kernel methods [22, 23] or linear quantum models [25].
Code Availability
The source code for the numerical case studies is publicly accessible on GitHub via
@@ 6 Conclusion
https://github.com/daniel-fink-de/training-robust-and-generalizable-quantum-models.
Acknowledgment
@@
This work was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 - 390740016. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). This work was also supported by the German Federal Ministry of Economic Affairs and Climate Action through the project AutoQML (grant no. 01MQ22002A).
Appendix A Lipschitz bounds of quantum models
@@
In this section, we study Lipschitz bounds of quantum models as in (4). We first derive a Lipschitz bound which is less tight than the one in (5) but can be shown using a simple concatenation argument (Section A.1). Next, in Section A.2, we prove that (5) is indeed a Lipschitz bound.
A.1 Simple Lipschitz bound based on concatenation
@@
Theorem A.1.
The following is a Lipschitz bound of $f_{\Theta}$:
@@
$\lVert f(x)-f(y)\rVert$ (18)
$\leq$ $L_{1}\lVert f_{2}\circ\dots\circ f_{h}(x)-f_{2}\circ\dots\circ f_{h}(y)\rVert$
$\leq$ $\dots\leq L_{1}L_{2}\cdots L_{h}\lVert x-y\rVert.$
@@
Lipschitz bound of $g_{\mathrm{unitary}}$: It follows from [31, Theorem 2.2] that $L_{\mathrm{unitary}}=\sum_{j=1}^{N}\lVert H_{j}\rVert$ is a Lipschitz bound of $g_{\mathrm{unitary}}$.
@@
A.2 Proof that (5) is a Lipschitz bound
@@
$\mathrm{d}U_{\Theta}(x)$ (26)
$=$ $(\mathrm{d}U_{N,\Theta_{N}}(x))U_{N-1,\Theta_{N-1}}(x)\cdots U_{1,\Theta_{1}}(x)$
@@
$\mathrm{d}U_{\Theta}(x)(v)$ (31)
$=$ $-i\left(H_{N}U_{\Theta}(x)w_{N}^{\top}v+\dots+U_{\Theta}(x)H_{1}w_{1}^{\top}v\right).$
@@
$J_{\Theta}(x)$ (32)
$=$ $-i\left(H_{N}U_{\Theta}(x)\ket{0}w_{N}^{\top}+\dots+U_{\Theta}(x)H_{1}\ket{0}w_{1}^{\top}\right).$
@@
Thus, $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ is a Lipschitz bound of $U_{\Theta}(x)\ket{0}$ [59, p. 356].
@@
Appendix B Full version and proof of Theorem 4.1
We first state the main result in a general supervised learning setup, before applying it to the quantum model (4) considered in the paper. Consider a supervised learning setup with data samples $\{x_{k},y_{k}\}_{k=1}^{n}$ drawn independently and identically distributed from $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ according to some probability distribution $P$. We define the $\epsilon$-covering number of $\mathcal{Z}$ as follows.
Definition B.1.
(adapted from [7, Definition 1]) We say that $\hat{\mathcal{Z}}$ is an $\epsilon$-cover of $\mathcal{Z}$, if, for all $z\in\mathcal{Z}$, there exists $\hat{z}\in\hat{\mathcal{Z}}$ such that $\lVert z-\hat{z}\rVert\leq\epsilon$. The $\epsilon$-covering number of $\mathcal{Z}$ is
@@
    $\mathcal{Z}$ is compact such that the value $M\coloneqq\sup_{y,y^{\prime}\in\mathcal{Y}}\ell(y,y^{\prime})$ is finite, and
@@
For the following proof, we invoke the concept of $(K,\epsilon)$-robustness (adapted from [7, Definition 2]): The classifier $f$ is $(K,\epsilon)$-robust for $K\in\mathbb{N}$ and $\epsilon\in\mathbb{R}$, if $\mathcal{Z}$ can be partitioned into $K$ disjoint sets, denoted by $\{C_{i}\}_{i=1}^{K}$, such that the following holds: For all $k=1,\dots,n$, $(x,y)\in\mathcal{Z}$, $i=1,\dots,K$, if $(x_{k},y_{k}),(x,y)\in C_{i}$, then
@@
where we use the Lipschitz bound $L_{\ell}$ of $\ell$, the triangle inequality, and the Lipschitz bound $L_{f}$ of $f$, respectively. Using [7, Theorem 6], we infer that $f$ is $(\mathcal{N}(\frac{\gamma}{2},\mathcal{Z}),L_{\ell}\max\{1,L_{f}\}\gamma)$-robust for all $\gamma>0$. It now follows from [7, Theorem 1] that, for any $\delta>0$, with probability at least $1-\delta$ inequality (37) holds. ∎
Let us now combine the Lipschitz bound (5) and Lemma B.1 to state a tailored generalization bound for the considered class of quantum models $f_{\Theta}$, thus proving Theorem 4.1.
Theorem B.1.
Suppose
@@
    $\mathcal{Z}$ is compact such that the value $M\coloneqq\sup_{y,y^{\prime}\in\mathcal{Y}}\ell(y,y^{\prime})$ is finite.
@@
$|\mathcal{R}(f)-\mathcal{R}_{n}(f)|$ (40)
$\leq$ $\gamma L_{\ell}\max\Big{\{}1,2\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert\Big{\}}$
$+M\sqrt{\frac{2\mathcal{N}(\frac{\gamma}{2},\mathcal{Z})\ln 2+2\ln(\frac{1}{\delta})}{n}}.$
Theorem B.1 shows that the size of the data encoding and of the observable directly influences the generalization performance of the quantum model $f_{\Theta}$. In particular, for smaller values of $\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert$ and $\lVert\mathcal{M}\rVert$, the expected loss is closer to the empirical loss. The right-hand side of (40) contains two terms: The first one depends on the parameters of the quantum model $f_{\Theta}$ and characterizes its robustness via the derived Lipschitz bound, whereas the second term decays with increasing data length $n$. While (37) holds for arbitrary values of $\gamma>0$, it is not immediate which value of $\gamma$ leads to the smallest possible bound: Smaller values of $\gamma$ decrease the first term but increase $\mathcal{N}(\frac{\gamma}{2},\mathcal{Z})$ in the second term (and vice versa).
In contrast to existing QML generalization bounds [24, 45, 46, 47, 48], Theorem B.1 explicitly highlights the role of the model parameters via the Lipschitz bound. Using the additional flexibility of the parameter $\gamma$, it can be shown that the generalization bound (40) converges to zero when the data length $n$ approaches infinity. To this end, we use [60, Lemma 6.27] to upper bound the covering number
@@
$|\mathcal{R}(f)-\mathcal{R}_{n}(f)|$ (42)
$\leq$ $\frac{1}{n^{\frac{1}{2d+2}}}L_{\ell}\max\Big{\{}1,2\lVert\mathcal{M}\rVert\sum_{j=1}^{N}\lVert w_{j}\rVert\lVert H_{j}\rVert\Big{\}}$
$+M\sqrt{\frac{2\ln 2\left(6R\right)^{d+1}}{\sqrt{n}}+\frac{2\ln(\frac{1}{\delta})}{n}},$
@@
Appendix C Numerics: setup and further results
In the following, we provide details regarding the setup of our numerical results (Section C.1) and we present further numerical results regarding parameter regularization in quantum models with fixed encoding (Section C.2).
C.1 Numerical setup
@@
All numerical simulations within this work were performed using the Python QML library PennyLane [61]. As device, we used the noiseless simulator "lightning.qubit" together with the adjoint differentiation method, to enable fast and memory efficient gradient computations. In order to solve the optimization problem in (9), we apply the ADAM optimizer using a learning rate of $\eta=0.1$ and the suggested values for all other hyperparameters [62]. Furthermore, we run $200$ epochs throughout and train $12$ models based on different initial parameters for varying regularization parameters $\lambda\geq 0$. Adding the regularization does not introduce significant computational overhead as the evaluation of the cost only involves a weighted sum of the terms $\|w_{j}\|^{2}$. As final model, we take the set of parameters for the model with minimal cost over all runs and epochs. Furthermore, the training as well as the robustness and generalization analysis were parallelized using Dask [63].
@@
$\>U_{\mathrm{Rot}}(\alpha,\beta,\gamma)=R_{Z}(\gamma)R_{Y}(\beta)R_{Z}(\alpha)\,,$ (46)
$=\begin{pmatrix}e^{-\frac{i}{2}(\gamma+\alpha)}\cos(\frac{\beta}{2})&-e^{\frac{i}{2}(\gamma-\alpha)}\sin(\frac{\beta}{2})\\ e^{-\frac{i}{2}(\gamma-\alpha)}\sin(\frac{\beta}{2})&e^{\frac{i}{2}(\gamma+\alpha)}\cos(\frac{\beta}{2})\end{pmatrix}\,.$
Figure 6: Results for the generalization simulations for the circle classification problem when using the fixed-encoding quantum model (14) and regularization of the parameters $\phi$. The training and test setup are identical to simulations shown in Figures 2 and 3. We plot the dependency of the test accuracy and the Lipschitz bound on the hyperparameter $\lambda$ entering the regularized training problem (47). Further, we plot the test accuracy for the best solution of the unregularized training problem.
@@
C.2 Regularization in quantum models with fixed encoding
In the main text, we have seen that the Lipschitz bound of the quantum model with fixed encoding $f_{\phi}^{\mathrm{f}}(x)$ in (14) cannot be adapted by changing the parameters $\phi$. As a result, it is not possible to use Lipschitz bound regularization for improving robustness and generalization. In the following, we investigate whether regularization of $\phi$ can instead be used to improve generalization performance. More precisely, we consider the same numerical setup as for our generalization results with trainable encoding depicted in Figure 3. The main difference is that we consider a fixed-encoding quantum model $f_{\phi}^{\mathrm{f}}(x)$ (compare Figure 4) which is trained via the following regularized training problem
$\min_{\phi}\>\>\frac{1}{n}\sum_{k=1}^{n}\ell(f_{\phi}^{\mathrm{f}}(x_{k}),y_{k})+\lambda\sum_{j=1}^{L}\lVert\phi_{j}\rVert^{2}.$ (47)
The regularization with hyperparameter $\lambda>0$ aims at keeping the norms of the angles $\phi_{j}$ small. Figure 6 depicts the test accuracy and Lipschitz bound of the resulting quantum model for different regularization parameters $\lambda$. First, note that the Lipschitz bound is indeed constant for all choices of $\lambda$ due to the fixed encoding. Comparing Figures 3 and 6, we see that the trainable encoding yields a significantly higher test accuracy in comparison to the fixed encoding. Moreover, the influence of the regularization parameter on the test accuracy is much less pronounced for the fixed encoding than for the trainable encoding, confirming our previous discussion on the benefits of trainable encodings. The test accuracy is not entirely independent of $\lambda$ since 1) regularization of the parameters $\phi_{j}$ influences the optimization and can improve or deteriorate convergence and 2) biasing $\phi$ towards zero may be beneficial if the underlying ground-truth distribution is better approximated by a quantum model with small values $\phi$. Whether 2) brings practical benefits is, however, highly problem-specific as it depends on the distribution generating the data and on the circuit ansatz.